# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ Abstract
This paper develops structure-preserving, oscillation-eliminating discontinuous Galerkin (OEDG) schemes for ideal magnetohydrodynamics (MHD), as a sequel to our recent work [1]. The schemes are based on a locally divergence-free (LDF) oscillation-eliminating (OE) procedure to suppress spurious oscillations while maintaining many of the good properties of original DG schemes, such as conservation, local compactness, and optimal convergence rates. The OE procedure is built on a simple damping equation—a linear ordinary differential equation (ODE) whose solution can be exactly formulated. Because this OE procedure does not interfere with DG spatial discretization and RK stage update, it can be easily incorporated to existing DG codes as an independent module. These features make the proposed LDF OEDG schemes highly efficient and easy to implement. In addition, we present a positivity-preserving (PP) analysis of the LDF OEDG schemes on Cartesian meshes via the optimal convex decomposition technique and the geometric quasi-linearization (GQL) approach. Efficient PP LDF OEDG schemes are obtained with the HLL flux under a condition accessible by the simple local scaling PP limiter. Several one- and two-dimensional MHD tests confirm the accuracy, effectiveness, and robustness of the proposed structure-preserving OEDG schemes.
@@ 1 Introduction
The magnetohydrodynamic (MHD) equations are widely utilized for modeling macro-scale behaviors in various domains, including laboratory, space, and astrophysical plasmas. Notable applications of these equations can be found in studies of phenomena such as coronal mass ejections [2], the evolution of solar coronal magnetic fields [3], and changes within Earth’s magnetosphere [4]. As a nonlinear hyperbolic system, the ideal compressible MHD equations pose significant challenges in obtaining analytical solutions due to their complexity. Consequently, numerical methods have become the principal approach for deriving solutions in this field. Among these, discontinuous Galerkin (DG) methods are particularly favored for their capacity to achieve high-order accuracy, excellent local compactness, and compatibility with arbitrary mesh configurations (see, e.g., [5, 6] and references therein). These attributes have led to their widespread adoption in MHD numerical simulations, as evidenced by various studies (e.g., [7, 8, 9, 10, 11, 12]). However, when addressing MHD problems featuring strong discontinuities, such as shocks, conventional high-order DG methods can produce spurious oscillations near these discontinuities, potentially leading to non-physical outcomes and simulation failures.
@@
To suppress spurious oscillations in the DG methods for hyperbolic conservation laws, various techniques have been developed, including the total variation bounded limiter [5], the weighted essentially non-oscillatory limiter [13], and artificial diffusion [14]. These strategies effectively control oscillations and enhance the stability of DG schemes. However, they may compromise certain desirable attributes of the original DG schemes (e.g., optimal error estimates or superconvergence) or depend on some problem-specific parameters. Recently, Lu, Liu, and Shu [15] proposed the so-called oscillation-free DG (OFDG) method for scalar conservation laws, which skillfully integrates a damping term into the semi-discrete DG schemes. The OFDG method can mitigate numerical oscillations near strong discontinuities while preserving many advantageous properties of the conventional DG schemes, including conservation, superconvergence, and optimal error estimates. Subsequently, the OFDG method was successfully extended to hyperbolic systems [16] with characteristic decomposition and applied to shallow water equations [17], nonlinear degenerate parabolic equations [18], and multi-component chemically reacting flows [19]. However, the characteristic decomposition in the OFDG method may be computationally expensive and require careful adjustments to the eigenvectors due to the lack of scale invariance. Furthermore, as Liu, Lu, and Shu have mentioned [16], the stiffness of the large damping terms near strong discontinuities makes the Courant–Friedrichs–Lewy (CFL) condition more restrictive, necessitating the use of modified exponential Runge–Kutta methods to alleviate this restriction. More recently, motivated by [15], a novel oscillation-eliminating DG (OEDG) method was systematically proposed in [1] by alternately evolving the conventional semi-discrete DG scheme and a new damping equation. Compared to the OFDG method, the OEDG method decouples the damping terms from the DG formulations, thereby being non-intrusive for integration into existing DG codes and simpler and more efficient for implementation. Moreover, the OEDG method is free from characteristic decomposition for hyperbolic systems, is stable under the normal CFL condition without requiring the exponential Runge–Kutta discretization, and exhibits desirable scale-invariant and evolution-invariant properties for problems spanning various scales and wave speeds.
@@
In ideal MHD, the magnetic field is divergence-free, and both density and internal energy (or pressure) are positive. Maintaining these conditions—zero divergence of the magnetic field and positivity of density and pressure—poses significant challenges in MHD simulations. Large divergence errors can induce nonphysical features or lead to numerical instabilities [20, 21]. To reduce the divergence errors, various techniques have been developed, including the projection method [20], constrained transport method [22, 23], eight-wave method with the Godunov–Powell source term [24], hyperbolic divergence cleaning method [25], locally divergence-free (LDF) method [7, 26], and globally divergence-free (GDF) method [27, 9, 28, 29, 12]. These approaches have been extensively applied across numerous schemes, thanks to their efficacy in controlling divergence. Positivity of density and pressure can be compromised in challenging MHD simulations involving conditions such as low density or pressure, high Mach numbers, large magnetic energies, or strong discontinuities. To address this, early research primarily focused on developing one-dimensional robust approximate Riemann solvers; see, e.g., [30, 31, 32, 33]. Waagan [34] introduced a second-order MUSCL-Hancock scheme using relaxation Riemann solvers [32, 33] that ensures the positivity of density and pressure; the robustness of this scheme was further validated through extensive testing [35]. There has been notable progress in the development of positivity-preserving (PP) limiters for higher-order schemes [36, 37]. Balsara [38] proposed a self-adjusting PP limiter to maintain the positivity of reconstructed density and pressure in ideal MHD. Cheng et al. [39] extended the PP limiter from [37] and showed the PP property of one-dimensional (central) DG schemes under some assumptions for ideal MHD. Christlieb et al. [40, 41] developed high-order PP finite difference methods for ideal MHD utilizing a parameterized flux limiter. These PP techniques have enhanced the robustness of MHD schemes, as evidenced by numerous numerical tests. However, the intrinsic complexity of MHD equations and the unclear relationship between the PP property and the divergence-free condition of the magnetic field have limited the proposal of numerical schemes that can rigorously prove to preserve positivity for MHD in earlier research.
@@
In recent years, significant advances [42, 43, 44, 45, 46, 47] have been made in revealing the theoretical connections between the PP property and the divergence-free condition for MHD systems. The work [42] first rigorously proved that a discrete divergence-free (DDF) condition is intrinsically linked to achieving the PP property for the ideal MHD equations. As discovered in [42], the PP property may be compromised even if there is a minor violation of the DDF condition. Furthermore, it has been found in [43] that even the exact smooth solutions of the conservative MHD system with nonzero magnetic divergence may result in negative pressure. The symmetrizable MHD system (i.e., the MHD system with the Godunov–Powell source term [48]) does not exhibit this problem [44], and a provably PP DG scheme for this system was first developed in [43]. Subsequently, [44] established a framework for provably positive high-order finite volume and DG schemes on general meshes. All these findings are based on an innovative technique known as geometric quasilinearization (GQL), which simplifies nonlinear constraints—such as the positivity of internal energy (or pressure) in the MHD system—into linear constraints by incorporating suitable auxiliary variables. Building on the geometric properties of convex sets, the general GQL framework was recently proposed in [45] to study bound-preserving schemes under nonlinear constraints. For the related developments in relativistic MHD, see [49, 50, 51]. Very recently, a second-order structure-preserving finite element method was proposed in [52] based on convex limiting and a novel splitting.
@@
In this paper, we develop the LDF OEDG method for the ideal MHD and carry out a PP analysis of the OEDG method with the HLL flux. The contributions, innovations, and significance of this work are outlined as follows:
@@
    For extending the OEDG method to MHD, how to make the OEDG method and the LDF condition compatible is something that must be considered. In order to control the oscillations without losing the LDF property of the computed magnetic field, we perform OE steps for Legendre polynomials and LDF polynomials respectively using the LDF DG method as the base scheme, and propose an LDF OEDG scheme. The LDF OEDG scheme can inherently balance LDF constraint and oscillation suppression, and is computationally cheap for arbitrary meshes.
@@
    In the framework that proves PP property, it is well known that the decomposition of the cell averages of the numerical solution into a convex combination of the solution values at certain quadrature points is a key ingredient. By employing the LDF OEDG scheme, PP limiter, HLL flux and properly discretized Godunov-Powell source term, we provide a provably PP analysis via general convex decomposition [53] on Cartesian meshes, and also give derived CFL conditions via some specific convex decompositions, including Zhang-Shu convex decomposition [36] and optimal convex decomposition [54].
@@
    In addition, different from the works in [44], we only consider adding Godunov-Powell source term to the evolution equations of the cell averages, since the PP property of the scheme is only related to the discrete equations satisfied by the cell averages.
Some one- and two-dimensional (2D) MHD examples are utilized to test the accuracy, effectiveness, and robustness of the proposed scheme. The numerical results show that our scheme can achieve expected convergence order, effectively suppress oscillations, and behaves robust in some tough problems.
The paper is organized as follows. In section 2, we proposed a LDF OEDG method for 2D ideal MHD equations. In section 3, we construct a PP LDF OEDG scheme for 2D ideal MHD equations and prove the PP property of the scheme. In section 4, we show several numerical examples for MHD in one and two dimensions. Section 5 gives the concluding remarks for this paper.
@@ 2 Locally divergence-free OEDG method for ideal MHD
This section presents the LDF OEDG method for the ideal MHD system. Without loss of generality, we focus on the 2D MHD equations, while our method is directly applicable to the one-dimensional case and can be readily extended to the three-dimensional case.
@@ 2.1 Governing equations of 2D MHD system
where $\boldsymbol{U}=\left(\rho,\rho\boldsymbol{u},\boldsymbol{B},E\right)^{\top}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}$ is the conservation vector, and the fluxes in the $x$- and $y$-directions are respectively given by
$\boldsymbol{F}_{1}\left(\boldsymbol{U}\right)=\left(\begin{array}[]{c}\rho u_{1}\\ \rho{u_{1}}^{2}+\left(p+{\|\boldsymbol{B}\|}^{2}/2\right)-B_{1}^{2}\\ \rho u_{1}u_{2}-B_{1}B_{2}\\ \rho u_{1}u_{3}-B_{1}B_{3}\\ 0\\ u_{1}B_{2}-B_{1}u_{2}\\ u_{1}B_{3}-B_{1}u_{3}\\ \left(E+p+{\|\boldsymbol{B}\|}^{2}/2\right)u_{1}-B_{1}\left(\boldsymbol{B}\cdot\boldsymbol{u}\right)\\ \end{array}\right),\quad\boldsymbol{F}_{2}\left(\boldsymbol{U}\right)=\left(\begin{array}[]{c}\rho u_{2}\\ \rho u_{2}u_{1}-B_{2}B_{1}\\ \rho{u_{2}}^{2}+\left(p+{\|\boldsymbol{B}\|}^{2}/2\right)-B_{2}^{2}\\ \rho u_{2}u_{3}-B_{2}B_{3}\\ u_{2}B_{1}-B_{2}u_{1}\\ 0\\ u_{2}B_{3}-B_{2}u_{3}\\ \left(E+p+{\|\boldsymbol{B}\|}^{2}/2\right)u_{2}-B_{2}\left(\boldsymbol{B}\cdot\boldsymbol{u}\right)\\ \end{array}\right).$
@@
$\nabla\cdot\boldsymbol{B}=0,$ (2)
@@
and correspondingly divide the flux $\boldsymbol{F}{i}(\boldsymbol{U})$ into $\boldsymbol{F}{i}^{R}(\boldsymbol{U})$ and $\boldsymbol{F}{i}^{Q}(\boldsymbol{U})$. The MHD system (1) is then split into

### v2

@@ Abstract
Numerically simulating magnetohydrodynamics (MHD) poses notable challenges, including the suppression of spurious oscillations near discontinuities (e.g., shocks) and preservation of essential physical structures (e.g., the divergence-free constraint of magnetic field and the positivity of density and pressure). This paper develops structure-preserving oscillation-eliminating discontinuous Galerkin (OEDG) schemes for ideal MHD, as a sequel to our recent work [M. Peng, Z. Sun & K. Wu, arXiv:2310.04807, 2023]. The schemes leverage a locally divergence-free (LDF) oscillation-eliminating (OE) procedure to suppress spurious oscillations while retaining the LDF property of magnetic field and many desirable attributes of original DG schemes, such as conservation, local compactness, and optimal convergence rates. The OE procedure is based on the solution operator of a novel damping equation, a linear system of ordinary differential equations that are exactly solvable without any discretization. The OE procedure is performed after each Runge–Kutta stage and does not impact DG spatial discretization, facilitating its easy integration into existing DG codes as an independent module. Moreover, this paper presents a rigorous positivity-preserving (PP) analysis of the LDF OEDG schemes on Cartesian meshes, utilizing the optimal convex decomposition technique [S. Cui, S. Ding & K. Wu, SIAM J. Numer. Anal., 62:775–810, 2024] and the geometric quasi-linearization (GQL) approach [K. Wu & C.-W. Shu, SIAM Review, 65:1031–1073, 2023]. Efficient PP LDF OEDG schemes are derived by incorporating appropriate discretization of Godunov–Powell source terms into only the discrete equations of cell averages, under a condition achievable through a simple PP limiter. Several one- and two-dimensional MHD tests verify the accuracy, effectiveness, and robustness of the proposed structure-preserving OEDG schemes.
@@ 1 Introduction
The magnetohydrodynamic (MHD) equations are extensively employed to model macro-scale behaviors across various domains, such as laboratory, space, and astrophysical plasmas. These equations find notable applications in the study of phenomena including coronal mass ejections [1], the evolution of solar coronal magnetic fields [2], and the dynamics within Earth’s magnetosphere [3]. As a nonlinear hyperbolic system, the ideal compressible MHD equations are inherently complex, making analytical solutions challenging to derive. Thus, numerical methods have become a primary approach in this research area. Among these methods, discontinuous Galerkin (DG) methods are especially preferred due to their ability to achieve high-order accuracy, excellent local compactness, and adaptability to arbitrary mesh configurations (cf. [4, 5]). The effectiveness of DG methods in MHD numerical simulations is well-documented in various studies (e.g., [6, 7, 8, 9, 10, 11]). However, in scenarios with strong discontinuities like shocks, conventional high-order DG methods may generate spurious oscillations near these discontinuities, leading to non-physical results and potential simulation failures.
@@
To suppress spurious oscillations in the DG methods for hyperbolic conservation laws, various techniques have been developed, including but not limited to the total variation bounded limiter [4], the weighted essentially non-oscillatory limiter [12], and artificial viscosity techniques (e.g., [13, 14]). These strategies effectively control oscillations and enhance the stability of DG schemes. However, some of them may compromise certain desirable attributes of the original DG schemes, such as optimal convergence rates, or depend on problem-specific parameters. Recently, Lu, Liu, and Shu [15] introduced the so-called oscillation-free DG (OFDG) method for scalar conservation laws, by ingeniously incorporating a damping term into semi-discrete DG schemes. The OFDG method can effectively mitigate numerical oscillations near strong discontinuities while retaining many beneficial features of the original DG schemes, including conservation, superconvergence, and optimal error estimates. Subsequently, the OFDG method was successfully extended to hyperbolic systems [16] with characteristic decomposition and applied to shallow water equations [17], nonlinear degenerate parabolic equations [18], and multi-component chemically reacting flows [19]. However, the characteristic decomposition in the OFDG method may be computationally expensive and require careful adjustments to the eigenvectors due to a lack of scale invariance. Furthermore, as Liu, Lu, and Shu mentioned [16], the stiffness of the large damping terms near strong discontinuities makes the Courant–Friedrichs–Lewy (CFL) condition more restrictive, necessitating the use of modified exponential Runge–Kutta methods to alleviate this restriction. More recently, motivated by [15], a novel oscillation-eliminating DG (OEDG) method was systematically developed in [20]. This method alternately evolves the conventional semi-discrete DG scheme and a novel damping equation. Unlike the OFDG method, the OEDG approach decouples the damping terms from the DG formulations, facilitating non-intrusive integration into existing DG codes, simplifying implementation, and enhancing efficiency. Additionally, the OEDG method does not require characteristic decomposition for hyperbolic systems, maintains stability under the standard CFL condition without the need for exponential Runge–Kutta discretization, and exhibits desirable properties such as scale invariance and evolution invariance across various scales and wave speeds.
@@
In ideal MHD, the magnetic field must be divergence-free, and both density and internal energy (or pressure) must remain positive. Preserving these physical structures—zero divergence of the magnetic field and positivity of density and pressure—presents significant challenges in MHD simulations. Large divergence errors can introduce nonphysical features or cause numerical instabilities [21, 22]. To mitigate these divergence errors, several techniques have been developed, including but not limited to the projection method [21], the constrained transport method [23, 24], the eight-wave method [25], the hyperbolic divergence cleaning method [26], the locally divergence-free (LDF) method [6, 27, 28, 29], and the globally divergence-free (GDF) method [30, 7, 31, 32, 11]. These methods have been widely applied across various schemes due to their effectiveness in controlling divergence errors. Positivity of density and pressure can be compromised in challenging MHD simulations involving low density or pressure, high Mach numbers, large magnetic energies, or strong discontinuities. To address this, early research primarily focused on developing robust one-dimensional approximate Riemann solvers; see, for example, [33, 34, 35, 36]. Waagan [37] introduced a second-order MUSCL-Hancock scheme using relaxation Riemann solvers [35, 36] that ensures the positivity of density and pressure; the robustness of this scheme was further validated through extensive testing [38]. There has been notable progress in the development of positivity-preserving (PP) limiters for higher-order schemes [39, 40]. Balsara [41] proposed a self-adjusting PP limiter to maintain the positivity of reconstructed density and pressure in ideal MHD. Cheng et al. [42] extended the PP limiter of Zhang–Shu [40] to ideal MHD and demonstrated the PP property of one-dimensional (central) DG schemes under certain assumptions. Christlieb et al. [43, 44] developed high-order PP finite difference methods for ideal MHD, via a parameterized flux limiter [45]. These PP techniques have significantly enhanced the robustness of MHD schemes, as evidenced by numerous numerical tests. However, the inherent complexity of MHD equations and the unclear relationship between the PP property and the divergence-free condition have limited earlier research in proposing numerical schemes that can be rigorously proven to preserve positivity for ideal MHD.
@@
In recent years, important advances [46, 47, 48, 49, 50] have been made in revealing the theoretical connections between the PP property and the divergence-free condition for MHD systems. For the ideal MHD equations, the work [46] first rigorously proved that a discrete divergence-free condition is intrinsically linked to achieving the PP property. As discovered in [46], even slight deviations from the discrete divergence-free condition could compromise the PP property. Furthermore, it has been found in [47] that even the exact smooth solutions of the conservative MHD system with nonzero magnetic divergence may result in negative pressure. Fortunately, as shown in [48], this issue does not arise in the symmetrizable MHD system, which incorporates the Godunov–Powell source term [51, 52]. A provably PP DG scheme for this system was first developed in [47]. Subsequently, [48] established a framework for provably positive high-order finite volume and DG schemes on general meshes. These advancements leverage an innovative approach termed geometric quasilinearization (GQL) [46, 49], which equivalently transforms complex nonlinear constraints—such as ensuring the positivity of internal energy or pressure in the MHD system—into manageable linear constraints through the integration of appropriate auxiliary variables. Capitalizing on the geometric characteristics of convex sets, a comprehensive GQL framework was recently proposed in [49] to address bound-preserving schemes under nonlinear constraints. For developments in the field of relativistic MHD, refer to [53, 54, 55]. More recently, a novel second-order structure-preserving finite element method was introduced in [56], based on convex limiting and a novel splitting technique.
@@
This paper aims to develop the LDF OEDG method for ideal MHD and then present a rigorous PP analysis. Our analysis establishes the probable PP property for the LDF OEDG method through the optimal convex decomposition technique [57, 58] and the GQL approach [49]. The contributions, novelty, and significance of this work are outlined as follows:
@@
    We design an LDF OE procedure, which effectively eliminates spurious oscillations while locally maintaining the divergence-free magnetic field. This procedure is integrated as an independent module within the LDF DG framework, applied after each Runge–Kutta stage, resulting in the LDF OEDG method. The procedure is built on the evolution operator of a novel damping equation, whose exact solution is explicitly formulated without any discretization. It serves as an LDF filter to the DG modal coefficients and is easily implemented and efficient, involving merely the multiplication of modal coefficients by scalars. The LDF OEDG method retains the desirable features of original DG schemes, such as conservation, local compactness, and optimal convergence rates, and is naturally extensible to general meshes.
@@
    A rigorous PP analysis of the LDF OEDG schemes with the Harten–Lax–van Leer (HLL) flux is carried out on Cartesian meshes. This analysis employs the recently developed GQL approach [46, 49], which skillfully transforms the nonlinear constraint of pressure into a family of equivalent linear constraints by introducing several auxiliary variables. To ensure the weak PP property for the cell averages of the OEDG solution, we incorporate suitable upwind discretization of Godunov–Powell source terms into the schemes. This addresses the impact of divergence errors on the PP property. Unlike our previous work [47, 48], we establish a couple of “two-state” inequalities for estimating the effect of the HLL flux on PP on rectangular cells, which simplifies/improves the estimates in [48] for general meshes when reduced to Cartesian meshes. Another difference is that we incorporate the discrete Godunov–Powell source terms solely into the evolution equations of the cell averages, as the weak PP property relates only to these discrete equations, thereby enhancing simplicity and efficiency.
@@
    A key aspect of the PP analysis is the convex decomposition, where the cell averages of the numerical solution are decomposed into a convex combination of its point values at certain quadrature points. We provide a comprehensive PP analysis via general convex decomposition [58] on Cartesian meshes and derive PP CFL conditions through specific convex decompositions, including the Zhang–Shu convex decomposition [39] and the optimal convex decomposition [57, 58]. The optimal convex decomposition leads to a milder theoretical PP CFL condition and requires fewer nodes, thus improving efficiency further. Our theoretical analyses establish the weak PP property of the LDF OEDG method, implying that employing a simple scaling PP limiter can enforce the pointwise PP property of the OEDG solution at any point of interest.
Some one- and two-dimensional (2D) MHD examples are utilized to test the accuracy, effectiveness, and robustness of the proposed PP LDF OEDG method. The numerical results demonstrate that our method achieves the expected optimal convergence order, effectively suppresses oscillations, and performs robustly in challenging scenarios.
The paper is organized as follows. Section 2 presents the LDF OEDG method for 2D ideal MHD equations. In section 3, we construct a PP LDF OEDG scheme for 2D ideal MHD equations and prove the PP property. Section 4 provides several numerical examples for MHD in one and two dimensions. Section 5 gives the concluding remarks of this paper.
@@ 2 Locally divergence-free OEDG method for ideal MHD
This section presents the LDF OEDG method for the ideal MHD system. Without loss of generality, we focus on the 2D MHD equations, while our method is directly applicable to the one-dimensional MHD system and can be readily extended to the three-dimensional case.
@@ 2.1 Governing equations of 2D MHD system
where $\boldsymbol{U}=\left(\rho,\rho\boldsymbol{u},\boldsymbol{B},E\right)^{\top}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}$ is the conservation vector. In alignment with existing studies in the literature, this paper considers the general configurations where $\boldsymbol{m},\boldsymbol{B}\in\mathbb{R}^{3}$ rather than $\mathbb{R}^{2}$, to accommodate the general 2D case that is a reduction from the 3D scenario. The fluxes in the $x$- and $y$-directions are respectively given by
$\boldsymbol{F}_{1}\left(\boldsymbol{U}\right)=\begin{pmatrix}\rho u_{1}\\ \rho{u_{1}}^{2}+\left(p+{\|\boldsymbol{B}\|}^{2}/2\right)-B_{1}^{2}\\ \rho u_{1}u_{2}-B_{1}B_{2}\\ \rho u_{1}u_{3}-B_{1}B_{3}\\ 0\\ u_{1}B_{2}-B_{1}u_{2}\\ u_{1}B_{3}-B_{1}u_{3}\\ \left(E+p+{\|\boldsymbol{B}\|}^{2}/2\right)u_{1}-B_{1}\left(\boldsymbol{B}\cdot\boldsymbol{u}\right)\\ \end{pmatrix},\quad\boldsymbol{F}_{2}\left(\boldsymbol{U}\right)=\begin{pmatrix}\rho u_{2}\\ \rho u_{2}u_{1}-B_{2}B_{1}\\ \rho{u_{2}}^{2}+\left(p+{\|\boldsymbol{B}\|}^{2}/2\right)-B_{2}^{2}\\ \rho u_{2}u_{3}-B_{2}B_{3}\\ u_{2}B_{1}-B_{2}u_{1}\\ 0\\ u_{2}B_{3}-B_{2}u_{3}\\ \left(E+p+{\|\boldsymbol{B}\|}^{2}/2\right)u_{2}-B_{2}\left(\boldsymbol{B}\cdot\boldsymbol{u}\right)\\ \end{pmatrix}.$
@@
$\nabla\cdot\boldsymbol{B}:=\frac{\partial B_{1}}{\partial x}+\frac{\partial B_{2}}{\partial y}=0,$ (2)
@@
and correspondingly divide the flux $\boldsymbol{F}_{i}(\boldsymbol{U})$ into $\boldsymbol{F}_{i}^{R}(\boldsymbol{U})$ and $\boldsymbol{F}_{i}^{Q}(\boldsymbol{U})$. The MHD system (1) is then split into